Example 12. Calculate the pH of a 1.0 x 10-8 M solution of HCI.Calculate the pH of a 1.0 x 10-8 M solution of HCl.Calculate the pH of a 0.0500 M solution of the weak acid nitrous acid (pKa = 3.5).
Answer #1

pH LH7 wCh ntt c ca9) PM-l09 C1ox108) eh=8 HNO 0.09 K-s0.0 a-16x16 nis very small Ka = OOS-7 s616X 101 (005) 11 ph= -3 (3.47